Hi, I have a word problem. And it says that I have to assume that the exercise describes a linear relationship. ok... here it is.
A vendor has learned that, by pricing hot dogs at $1.00, sales would reach 102 hot dogs per day. Raising the price to $1.50 will cause the sales to fall to 78 hot dogs per day. Let y be the number of hotdogs the vendor sells at x dollars each. Write a linear equation that models the number of hot dogs sold per day when the price is x dollars each.
In my work I keep getting 48x but Im not sure how. I may be negative, I don't know. And I'm getting both 54 and 150 to add on to it in the equation but again, I'm not sure. THANK YOU!

Answer by checkley71(8403)

THE TWO POINTS TO GRAPH THE PROBLEM ARE (100,102)(150,78) THUS THE SLOPE OF THE LINE IS (78-102)/(150-100) OR -24/50 OR -.48 THEN THE EQUATION FOR THIS LINE IS Y=-.48X+B NOW REPLACING Y & X BY ONE OF THE POINTS WE FIND THE VALUE OF B THUS 102=-.48*100+B OR 102=-48+B OR 102+48=B OR B=150 THUS THE EQUATION FOR THIS LINE/RELATIONSHIP IS Y=-.48X+150